OpenStudy (anonymous):

Robby invested $18,400 at 7% interest. He earned $9,016 interest at the end of x years. What is the value of x?

OpenStudy (anonymous):

Im wondering the same...

OpenStudy (anonymous):

I was having difficulty with this a few minutes a go since I was sick when the teacher taught this. Anyway, you would have to use the formula P*T*R=I P=Principle=18400 R=Rate=7%=0.07 T=The number of years=X I=Interest=9016 Solve for X

OpenStudy (anonymous):

Or in this case, T

OpenStudy (anonymous):

I think that how you would do it, give me a moment >.> Again, wasn't taught this

OpenStudy (anonymous):

Yeah, I was right. 18400*0.07=1288 9016/1288=X

OpenStudy (anonymous):

Do you get it?

OpenStudy (ezraweise):

whats the anser

OpenStudy (anonymous):

I think you can divide 9016 by 1288 yourself..